OpenStudy (anonymous):

56% as a decimal number.

OpenStudy (vishweshshrimali5):

56% = 56/100 = 0.56

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

Keep in mind that x% = x/100